In recent years many European nations have suffered from relatively high unemployment. The table below shows the unemployment rates (%) for 8 European nations.

European unemployment rates
Austria: 4.3 England: 5.1 Denmark: 5.8 Sweden: 6.4
Italy: 8.6 Germany: 9.1 Greece: 9.7 Spain: 11.2

(a) Find and interpret the median.
median =  (Use 2 decimals)

A. Approximately 50% of the data lie within 1 standard deviation of this value.

B. Approximately 50% of the data lie below this value.

C. This value is the average of the data set.

D. Approximately 25% of the data lie below this value.



(b) Find the first quartile (Q1) and the third quartile (Q3).

Q1 = (Use 2 decimals)
Q3 = (Use 2 decimals)


(c) Find and interpret the mean.
mean =  (Use 2 decimals)

A. Approximately 50% of the data lie below this value.     

B. Approximately 25% of the data lie below this value.

C. Approximately 50% of the data lie within 1 standard deviation of this value.

D. This value is the average of the data set.
Answer #1

Data is

4.3 5.1 5.8 6.4 8.6 9.1 9.7 11.2

a) Median is an average of middle two values.

Median = (6.4 + 8.6) / 2 = 7.5

Median = 7.5

B. Approximately 50% of the data lie below this value.​

b)

Q1 = 5.63

( Using excel =QUARTILE(Select array,1))

Q2 = 9.25

( Using excel =QUARTILE(Select array,3))

c) Mean is average of the data set values.

Mean = 60.2 / 8 = 7.53

Mean = 7.53

D. This value is the average of the data set.​